OpenStudy (anonymous):

How do you simplify indices? Eg, (4g^2)^3 / 2(32g^2)^2

OpenStudy (anonymous):

\[\frac{4g^2\times4g^2\times4g^2}{2\times32g^2\times32g^2} = \frac{2g^2}{64} = \frac{g^2}{32}\]

OpenStudy (anonymous):

I don't see how you went from \[4g^2 \times 4g^2 \times 4g^2 \to 2g^2\] And from \[2 \times 32g^2 \times 32g^2 \to 64\]

OpenStudy (anonymous):

I skipped some steps in between. Sorry about that \[\frac{4g^2\times4g^2\times4g^2}{2\times32g^2\times32g^2} = \frac{2g^2\times g^2\times g^2}{8g^2\times8g^2} = \frac{2g^2}{8\times8} = \frac{2g^2}{64}\]

OpenStudy (anonymous):

Cool, thanks :)

OpenStudy (anonymous):

No problems :) Hope it help